OpenStudy (anonymous):

2NH3+22 kcal=N2+ 6H how much heat energy must be absorbed during the production of 9 moles of hydrogen gas? please help, I'm stuck with where to start the equation!

OpenStudy (aaronq):

the equation, as is written, is incorrect. It should read: \(2NH_3+~22 kcal\rightleftharpoons N_2+ 6H_2\) now we need to use the coefficient for hydrogen gas, \(\sf Energy~absorbed=\dfrac{moles~of~H_2}{H_2's~coefficient}*E_{reaction}\) \(\sf Energy~absorbed=\dfrac{9~moles}{6}*22~kcal/mol=\dfrac{9*22}{6}~ kcal \)